OpenStudy (anonymous):

What is the equation of y equals x cubed with the given transformations? vertical stretch by a factor of 3, horizontal shift 4 units to the right, vertical shift 3 units down

OpenStudy (danjs):

Horizontal shift of f(x) ---> f(x + a) Vertical shif f(x) ----f(x) +a

OpenStudy (danjs):

stretch f(x) --->a*f(x)

OpenStudy (danjs):

Use f(x) = x^3 with those transformations

OpenStudy (kc_kennylau):

Horizontal stretch: \(\displaystyle y=f(\frac{x}{k})\) Vertical stretch: \(y=kf(x)\) Horizontal shift: \(y=f(x-k)\) Vertical shift: \(y=f(x)+k\) By the magnitude of \(k\).

OpenStudy (anonymous):

So I have y=3(x^3-4)+3?

OpenStudy (kc_kennylau):

Nope, perform the steps one by one

OpenStudy (kc_kennylau):

it should be (x-4)^3 instead of x^3-4, for that matter

OpenStudy (anonymous):

Okay so to start I do the vertical stretch by 3 y=3 (x^3) Then the horizontal y=3(x-4)^3 Then the vertical shift y=3(x-4)^3-3 Like this?

OpenStudy (kc_kennylau):

Yep, exactly.

OpenStudy (anonymous):

Thanks!